The perimeter of a rectangle is 120 feet. The width of the rectangle is equal to one-half the length and width of the rectangle
Answer by mananth(16949)

let the length of the rectangle be x
the width is 1/2 *x
the perimeter is 2*L + 2*w
120= 2*x+ 2* (1/2 *x)
120 = 2x+x
120 =3x
120/3= x
40 =x
the width will be 1/2 of 40 = 20 feet